OpenStudy (anonymous):

Bradley answered 80% of the question on his science test correctly. There were 30 questions on the test, and all the questions were equal value. How many questions did bradley NOT answer correctly

OpenStudy (anonymous):

6

OpenStudy (triciaal):

percent means per 100 80% correct means 100-80 = 20% incorrect total is 30 20% of 30 = (20/100)(30)

OpenStudy (***[isuru]***):

if she answers 80% correctly... the remaining 20% is the amount that she failed to answer correctly ... so find 20% of 30 (20/100) x 30 = 6%

OpenStudy (triciaal):

@***[ISURU]*** not % just 6 the amount of questions incorrect